# Reconciling $M/L$ Ratios Across Cosmic Time: a Concordance IMF for Massive Galaxies

### v1

@@ Abstract
The stellar initial mass function (IMF) is thought to be bottom-heavy in the cores of the most massive galaxies, with an excess of low mass stars compared to the Milky Way. However, studies of the kinematics of quiescent galaxies at $52<z<$ 5 find $M/L$ ratios that are better explained with lighter IMFs. Light IMFs have also been proposed for the unexpected populations of luminous galaxies that JWST has uncovered at $z>7$, to reduce tensions with galaxy formation models. Here we explore IMFs that are simultaneously bottom-heavy, with a steep slope at low stellar masses, and top-heavy, with a shallow slope at high masses. We derive a form of the IMF for massive galaxies that is consistent with measurements in the local universe and yet produces relatively low $M/L$ ratios at high redshift. This ‘concordance’ IMF has slopes $\gamma_{1}=2.40\pm 0.09$, $\gamma_{2}=2.00\pm 0.14$, and $\gamma_{3}=1.85\pm 0.11$ in the regimes 0.1 M – 0.5 M, 0.5 M – 1 M, and $>1$ M respectively. The IMF parameter $\alpha$, the mass excess compared to a Milky Way IMF, ranges from $\log(\alpha)\approx+0.3$ for present-day galaxies to $\log(\alpha)\approx-0.1$ for their star forming progenitors. The concordance IMF applies only to the central regions of the most massive galaxies, with velocity dispersions $\sigma\sim 300$ km s-1, and their progenitors. However, it can be generalized using a previously-measured relation between $\alpha$ and $\sigma$. We arrive at the following modification to the Kroupa (2001) IMF for galaxies with $\sigma\gtrsim 160$ km s-1: $\gamma_{1}\approx 1.3+4.3\log\sigma_{160}$; $\gamma_{2}\approx 2.3-1.2\log\sigma_{160}$; and $\gamma_{3}\approx 2.3-1.7\log\sigma_{160}$, with $\sigma_{160}=\sigma/160$ km s-1. If galaxies grow primarily inside-out, so that velocity dispersions are relatively stable, these relations should also hold at high redshift.
@@ 1 Introduction
The form of the stellar initial mass function (IMF), and its potential variation with star forming conditions, is a topic of ongoing debate (see the reviews by Bastian et al., 2010; Offner et al., 2014; Smith, 2020). The question has taken on a new significance since the launch of JWST, as unexpected populations of bright and seemingly massive galaxies have been discovered at $z>7$ (e.g., Naidu et al., 2022; Labbé et al., 2023; Xiao et al., 2023). The formation of the stellar component of these galaxies requires extremely rapid conversion of gas into stars, perhaps more rapid than can be accommodated in standard galaxy formation models (see Boylan-Kolchin, 2023; Dekel et al., 2023). There is a straightforward way to reduce or even eliminate this tension, as the total mass of a stellar population is dominated by stars that have low masses and contribute very little to the light. Therefore, as pointed out by Steinhardt et al. (2023) and others (Woodrum et al., 2023; Wang et al., 2024a), if the IMF in these galaxies were top-heavy (or bottom-light) their derived masses would be lowered considerably. Galaxy formation models with top-heavy IMFs also provide better fits to the UV luminosity function at these redshifts (Inayoshi et al., 2022; Harikane et al., 2023; Yung et al., 2024; Menon et al., 2024).
There is good evidence that such IMFs can occur in certain environments; as an example, star clusters in the central $\sim 300$ pc of the Milky Way appear to have an excess of high mass stars compared to young clusters elsewhere in the disk (Lu et al., 2013; Hosek et al., 2019; Chabrier & Dumond, 2024). Top-heavy IMFs for early galaxies have also been proposed on theoretical grounds, as higher CMB temperatures can lead to an increase in the turnover mass (Larson, 1998; Bate, 2023). However, the IMF in the central regions of massive early-type galaxies appears to be bottom-heavy, with an excess of low mass stars compared to the Milky Way (see Smith, 2020, for a review). The evidence for bottom-heavy IMFs comes primarily from studies of gravity-sensitive stellar absorption lines (Cenarro et al., 2003; van Dokkum & Conroy, 2010; Spiniello et al., 2012; Conroy & van Dokkum, 2012b; La Barbera et al., 2013; Martín-Navarro et al., 2015; La Barbera et al., 2019), which detect the low mass stars directly, and also from higher-than-expected stellar $M/L$ ratios derived from gravitational lensing (Treu et al., 2010; Posacki et al., 2015) and galaxy dynamics (Cappellari et al., 2012; Li et al., 2017; Shetty et al., 2020). As the old and dense central regions of massive galaxies in the present-day Universe are the likely descendants of the most luminous $z>7$ galaxies, these results suggest that – if anything – the masses of the JWST-discovered galaxies might be underestimated, by factors of 1.5–3.
@@ 2.1 General Functional Form
The IMFs are parameterized by the three power law segments that are familiar from the Kroupa (2001) IMF: 0.08 M$<m<0.5$ M, 0.5 M$<m<1$ M, and $m>1$ M. The power law slopes in these regions are denoted with $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$ respectively. The Kroupa (2001) IMF, appropriate for star formation in the Milky Way disk, has $\gamma_{1}=1.3$, $\gamma_{2}=2.3$, and $\gamma_{3}=2.3$. It is shown in Fig. 1, along with the (functionally nearly identical) Chabrier (2003) IMF and the Salpeter (1955) IMF (which has $\gamma_{1}=\gamma_{2}=\gamma_{3}=2.35$). We use integration limits of $0.1$ M – $100$ M.
@@
Figure 1: Illustration of the general form of the IMFs that are explored in this paper: a steep, Salpeter-like slope at low masses, followed by a flattening between $0.5$ M and $1$ M, and a further flattening at $>1$ M. Such IMFs can be described as being both bottom-heavy and top-heavy, or ‘middle-light’. For context and reference, the widely used Salpeter (1955), Kroupa (2001), and Chabrier (2003) forms are also shown.
@@ 2.2 Observational Constraints
Figure 2: Values of the IMF parameter $\alpha$ for all 484 ‘middle-light’ IMFs that are considered in this study, as a function of the three power law slopes $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$ that define the shape. In each panel the color palette is such that red is disfavored/disallowed and green is favored/allowed. Left: $\alpha_{\rm liv}$, the mass excess in living stars, for an age of 12.5 Gyr. Many IMFs satisfy the constraint $\alpha_{\rm liv}>1.6$. Middle: $\alpha_{\rm tot}$, the total mass excess, for an age of 12.5 Gyr. This is a strong function of the high mass slope, $\gamma_{3}$. Shallow high mass slopes produce too many remnants, and are disallowed by the criterion $\alpha_{\rm tot}<2.3$. Right: $\alpha_{\rm tot}$ for an age of 0.5 Gyr. The lowest values of $\alpha_{\rm tot}$ are in the range $1.1-1.3$.
@@ 3.2 Concordance IMF
where the uncertainty is taken to be the rms range among the 20 IMFs. This ‘concordance’ IMF is shown in Fig. 3. It is similar to the Salpeter IMF in terms of its ratio of low mass stars to turnoff stars, but flatter at high masses. The key turnover is at 0.5 M: the slope changes at this point in all 20 viable IMFs, by $0.4\pm 0.2$. The turnover at 1 M is not significant; of the 20 viable models, 7 have $\gamma_{2}=\gamma_{3}$.
@@ 4 Discussion
The physical origin of IMFs that are simultaneously bottom-heavy and top-heavy is unclear. The behavior of the IMF at low masses has long been linked to the Jeans mass and its dependence on temperature, density, and Mach number. In extreme environments the low-mass IMF can be supressed due to the $T^{3/2}$ temperature dependence of the Jeans mass (Larson, 1998; Krumholz, 2006; Bate, 2009, 2023), or enhanced if the $\rho^{-1/2}$ density dependence or Mach number dependence dominates (e.g., Hopkins, 2013; Chabrier et al., 2014; Tanvir et al., 2022; Tanvir & Krumholz, 2024), but it is difficult to do both at the same time (see, e.g., Bate, 2023). Models based on Press-Schecter-like arguments applied to turbulent molecular clouds (Hennebelle & Chabrier, 2008; Hopkins, 2013; Chabrier et al., 2014) directly link the high-mass IMF slope to the spectral index of the turbulent velocity power spectrum, with steeper spectra giving rise to shallower high-mass IMF slopes. Such models may provide a natural framework to simultaneously produce bottom-heavy and top-heavy IMFs by invoking turbulence as the key process working at both ends.
@@
It may also be that the observed IMF is a combination of star formation in different local environments within early galaxies, with the densest regions producing the steep low mass slope and the regions that are most exposed to radiation producing the flat high mass slope. JWST results have shown that many early galaxies have extremely high central surface densities (Baggen et al., 2023; Carnall et al., 2024), lending some observational support to this idea. Such hybrid IMFs have been developed previously in the framework of the ‘integrated galactic IMF’ (IGIMF; Kroupa & Weidner, 2003), and the IMF that Yan et al. (2021) and Haslbauer et al. (2024) propose for high metallicity, high star formation rate environments is qualitatively quite similar to the concordance IMF. Fontanot et al. (2018), working from similar concepts as Kroupa & Weidner (2003), introduce physically-motivated IMFs that have a break at 1 M with a steep slope at low masses and a shallow slope at high masses. As shown in Fontanot et al. (2023), these IMFs reproduce the $M/L$ ratios of present-day ellipticals well, although they underpredict the number of low mass stars (the mass excess with respect to the Milky Way IMF is largely in the form of remnants, not low mass stars, in these models).
@@
As a final note, we acknowledge the long history of appealing to IMF variation as a means of explaining real or imagined observational tensions, and that in such previous cases a more prosaic explanation was usually found. We therefore suggest adopting the concordance IMF with caution, for now.
@@
We thank Mariska Kriek and Bingjie Wang for their comments on the manuscript. Bingjie Wang also fitted the galaxies in her paper with the concordance IMF. We also thank the organizers of the 2011 Galaxy Formation conference in Durham, as this was when we first considered ‘ski slope’ IMFs as a potential way to reconcile seemingly contradictory results in the literature.

### v2

@@ Abstract
The stellar initial mass function (IMF) is thought to be bottom-heavy in the cores of the most massive galaxies, with an excess of low mass stars compared to the Milky Way. However, studies of the kinematics of quiescent galaxies at $52<z<$ 5 find $M/L$ ratios that indicate lighter IMFs. Light IMFs have also been proposed for the unexpected populations of luminous galaxies that JWST has uncovered at $z>7$, to reduce tensions with galaxy formation models. Here we explore ‘ski slope’ IMFs that are simultaneously bottom-heavy, with a steep slope at low stellar masses, and top-heavy, with a shallow slope at high masses. We derive a form of the IMF for massive galaxies that is consistent with measurements in the local universe and yet produces relatively low $M/L$ ratios at high redshift. This concordance IMF has slopes $\gamma_{1}=2.40\pm 0.09$, $\gamma_{2}=2.00\pm 0.14$, and $\gamma_{3}=1.85\pm 0.11$ in the regimes 0.08 M – 0.5 M, 0.5 M – 1 M, and $>1$ M respectively. The IMF parameter $\alpha$, the mass excess compared to a Milky Way IMF, ranges from $\log(\alpha)\approx+0.3$ for present-day galaxies to $\log(\alpha)\approx-0.1$ for their star forming progenitors. The concordance IMF applies only to the central regions of the most massive galaxies, with velocity dispersions $\sigma\sim 300$ km s-1, and their progenitors. However, it can be generalized using a previously-measured relation between $\alpha$ and $\sigma$. We arrive at the following modification to the Kroupa (2001) IMF for galaxies with $\sigma\gtrsim 160$ km s-1: $\gamma_{1}\approx 1.3+4.3\log\sigma_{160}$; $\gamma_{2}\approx 2.3-1.2\log\sigma_{160}$; and $\gamma_{3}\approx 2.3-1.7\log\sigma_{160}$, with $\sigma_{160}=\sigma/160$ km s-1. If galaxies grow primarily inside-out, so that velocity dispersions are relatively stable, these relations should also hold at high redshift.
@@ 1 Introduction
The form of the stellar initial mass function (IMF), and its potential variation with star forming conditions, is a topic of ongoing debate (see the reviews by Bastian et al., 2010; Offner et al., 2014; Smith, 2020). The question has taken on a new significance since the launch of JWST, as unexpected populations of bright and seemingly massive galaxies have been discovered at $z>7$ (e.g., Naidu et al., 2022; Labbé et al., 2023; Xiao et al., 2023). The formation of the stellar component of these galaxies requires extremely rapid conversion of gas into stars, perhaps more rapid than can be accommodated in standard galaxy formation models (see Boylan-Kolchin, 2023; Dekel et al., 2023). There is a straightforward way to reduce or even eliminate this tension, as the total mass of a stellar population is dominated by stars that have low masses and contribute very little to the light. Therefore, as pointed out by Steinhardt et al. (2023) and others (Woodrum et al., 2023; Wang et al., 2024a), if the IMF in these galaxies were top-heavy (or bottom-light) their derived masses would be lowered considerably. Galaxy formation models with top-heavy IMFs also provide better fits to the UV luminosity function at these redshifts (Inayoshi et al., 2022; Harikane et al., 2023; Yung et al., 2024; Schaerer et al., 2024).
There is good evidence that such IMFs can occur in certain environments; star clusters in the central $\sim 300$ pc of the Milky Way appear to have an excess of high mass stars compared to young clusters elsewhere in the disk (Lu et al., 2013; Hosek et al., 2019; Chabrier & Dumond, 2024), as does the 30 Doradus cluster in the Large Magellanic Cloud (Schneider et al., 2018). Top-heavy IMFs for early galaxies have also been proposed on theoretical grounds, as higher CMB temperatures can lead to an increase in the turnover mass (Larson, 1998; Bate, 2023). However, the IMF in the central regions of massive early-type galaxies appears to be bottom-heavy, with an excess of low mass stars compared to the Milky Way (see Smith, 2020, for a review). The evidence for bottom-heavy IMFs comes primarily from studies of gravity-sensitive stellar absorption lines (Cenarro et al., 2003; van Dokkum & Conroy, 2010; Spiniello et al., 2012; Conroy & van Dokkum, 2012b; La Barbera et al., 2013; Martín-Navarro et al., 2015; La Barbera et al., 2019), which detect the low mass stars directly, and also from higher-than-expected stellar $M/L$ ratios derived from gravitational lensing (Treu et al., 2010; Posacki et al., 2015) and galaxy dynamics (Cappellari et al., 2012; Li et al., 2017; Shetty et al., 2020). As the old and dense central regions of massive galaxies in the present-day Universe are the likely descendants of the most luminous $z>7$ galaxies, these results suggest that – if anything – the masses of the JWST-discovered galaxies might be underestimated, by factors of 1.5–3.
@@ 2.1 General Functional Form
The IMFs are parameterized by the three power law segments that are familiar from the Kroupa (2001) IMF: 0.08 M$<m<0.5$ M, 0.5 M$<m<1$ M, and $m>1$ M. The power law slopes in these regions are denoted with $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$ respectively. The Kroupa (2001) IMF, appropriate for star formation in the Milky Way disk, has $\gamma_{1}=1.3$, $\gamma_{2}=2.3$, and $\gamma_{3}=2.3$. It is shown in Fig. 1, along with the (functionally nearly identical) Chabrier (2003) IMF and the Salpeter (1955) IMF (which has $\gamma_{1}=\gamma_{2}=\gamma_{3}=2.35$). We use integration limits of $0.08$ M – $100$ M.
@@
Figure 1: Illustration of the general form of the IMFs that are explored in this paper: a steep, Salpeter-like slope at low masses, followed by a flattening between $0.5$ M and $1$ M, and a further flattening at $>1$ M. This gradual flattening somewhat resembles a ski slope. For context and reference, the widely used Salpeter (1955), Kroupa (2001), and Chabrier (2003) forms are also shown. Ski slope IMFs are simultaneously bottom-heavy and top-heavy compared to the IMF of the Milky Way.
@@ 2.2 Observational Constraints
Figure 2: Values of the IMF parameter $\alpha$ for all 484 ‘ski slope’ IMFs that are considered in this study, as a function of the three power law slopes $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$ that define the shape. In each panel the color palette is such that red is disfavored/disallowed and green is favored/allowed. Left: $\alpha_{\rm liv}$, the mass excess in living stars, for an age of 12.5 Gyr. Many IMFs satisfy the constraint $\alpha_{\rm liv}>1.6$. Middle: $\alpha_{\rm tot}$, the total mass excess, for an age of 12.5 Gyr. This is a strong function of the high mass slope, $\gamma_{3}$. Shallow high mass slopes produce too many remnants, and are disallowed by the criterion $\alpha_{\rm tot}<2.3$. Right: $\alpha_{\rm tot}$ for an age of 0.5 Gyr. The lowest values of $\alpha_{\rm tot}$ are in the range $1.1-1.3$.
@@ 3.2 Concordance IMF
where the uncertainty is taken to be the rms range among the 20 IMFs. This concordance IMF is shown in Fig. 3. It is similar to the Salpeter IMF in terms of its ratio of low mass stars to turnoff stars, but flatter at high masses. The key turnover is at 0.5 M: the slope changes at this point in all 20 viable IMFs, by $0.4\pm 0.2$. The turnover at 1 M is not significant; of the 20 viable models, 7 have $\gamma_{2}=\gamma_{3}$.
@@ 3.3 M/L Ratios Over Cosmic Time
3.4 Dependence on the Low Mass Cutoff
In our analysis the lower integration limit of the IMF is the hydrogen burning limit of 0.08 M (Chabrier et al., 2023), that is, it is assumed that the mass locked up in brown dwarfs is negligible. For a Milky Way IMF the precise lower limit is not important as most of the stellar mass is in stars near the turnover; for a Kroupa IMF, changing the lower limit to 0.03 M increases the total mass by 2 % and changing it to 0.01 M increases the mass by 3 %. However, for a Salpeter IMF the total mass is very sensitive to the low mass cutoff, increasing by a factor of 1.4 going from $0.08$ M to $0.03$ M. For a lower limit of 0.01 M the increase is a factor of 2.2, which means that most of the stellar mass is in the form of brown dwarfs.
As ski slope IMFs have a Salpeter-like low mass slope, their total mass is also very sensitive to the low mass cutoff. We repeated the analysis for integration limits of 0.03 M and 0.01 M, and find that no IMFs satisfy the critera of § 2.2. Relaxing the criteria to $\alpha_{\rm tot}<2.5$ (age = 12.5 Gyr) and $\alpha_{\rm tot}<1.5$ (age = 0.5 Gyr), there is one solution for a cutoff of 0.03 M, with $\gamma_{1}=2.3$, $\gamma_{2}=2.1$, and $\gamma_{3}=1.8$, and still no solution for a cutoff of 0.01 M.
We infer that a turnover near the hydrogen burning limit is required in the context of ski slope IMFs, as otherwise the total $M/L$ ratios would exceed constraints from dynamics and lensing. The steep slope at low masses and the sharp cutoff at $0.08$ M should be viewed as a simplified parameterization of a shift in the turnover mass from $\sim 0.3$ M for the Milky Way to $\sim 0.1$ M for the concordance IMF. The actual form of the low mass IMF may be similar to the IMFs explored in Chabrier et al. (2014), which combine slopes of $2.3-2.7$ with a turnover near $0.1$ M.
@@ 4 Discussion
The physical origin of IMFs that are simultaneously bottom-heavy and top-heavy is unclear. The behavior of the IMF at low masses has long been linked to the Jeans mass and its dependence on temperature, density, and Mach number. In extreme environments the low-mass IMF can be supressed due to the $T^{3/2}$ temperature dependence of the Jeans mass (Larson, 1998; Krumholz, 2006; Bate, 2009, 2023), or enhanced if the $\rho^{-1/2}$ density dependence or Mach number dependence dominates (e.g., Hopkins, 2013; Chabrier et al., 2014; Tanvir et al., 2022; Tanvir & Krumholz, 2024), but it is difficult to do both at the same time (see, e.g., Bate, 2023). A different process may operate at high masses. Models based on Press-Schecter-like arguments applied to turbulent molecular clouds (Hennebelle & Chabrier, 2008; Hopkins, 2013; Chabrier et al., 2014) directly link the high-mass IMF slope to the spectral index of the turbulent velocity power spectrum, with steeper spectra giving rise to shallower high-mass IMF slopes. These models have been tested and refined with hydrodynamical simulations (Nam et al., 2021), although it remains to be seen whether the required modifications to the spectral index occur in realistic environments. To summarize, $\gamma_{1}$ may be largely determined by the Jeans mass and $\gamma_{3}$ by the index of the turbulent power spectrum, but more work is needed to test this.
@@
It may also be that the observed IMF is a combination of star formation in different local environments within early galaxies, with the densest regions producing the steep low mass slope and the regions that are most exposed to radiation producing the flat high mass slope. JWST results have shown that many early galaxies have extremely high central surface densities (Baggen et al., 2023; Carnall et al., 2024), lending some observational support to this idea. Such hybrid IMFs have been developed previously in the framework of the ‘integrated galactic IMF’ (IGIMF; Kroupa & Weidner, 2003), and the IMF that Jeřábková et al. (2018), Yan et al. (2019, 2021), and Haslbauer et al. (2024) propose for high metallicity, high star formation rate environments is qualitatively quite similar to the concordance IMF. Fontanot et al. (2018), working from similar concepts as Kroupa & Weidner (2003), introduce physically-motivated IMFs that have a break at 1 M with a steep slope at low masses and a shallow slope at high masses. As shown in Fontanot et al. (2023), these IMFs reproduce the $M/L$ ratios of present-day ellipticals well, although they underpredict the number of low mass stars (the mass excess with respect to the Milky Way IMF is largely in the form of remnants, not low mass stars, in these models).
Several of the questions that are raised in the preceding paragraphs could be addressed if local analogs could be studied. Returning to the evidence for top-heavy IMFs discussed in the Introduction, the high mass slope in 30 Doradus is $1.90^{+0.37}_{-0.26}$ (Schneider et al., 2018), similar to that of the concordance IMF. It would be very interesting — although observationally challenging; see, e.g., Fahrion & De Marchi (2024) — to determine if there is an upturn at low masses in this cluster. We note that if such analogs exist in the Local Group, it would imply that the observed relations with global galaxy properties such as $\sigma$ are proxies for the preponderance of particular local conditions within the galaxies.
@@
As a final remark, we acknowledge the long history of appealing to IMF variation as a means of explaining real or imagined observational tensions, and that in such previous cases a more prosaic explanation was usually found. We therefore suggest adopting the concordance IMF with caution, for now.
@@
We thank Mariska Kriek and Bingjie Wang for their comments on the manuscript. Bingjie Wang also fitted the galaxies in her paper with the concordance IMF. We are grateful to the referee, Mark Krumholz, for insightful comments that improved the paper. Finally, we thank the organizers of the 2011 Galaxy Formation conference in Durham, as this was when we first considered ‘ski slope’ IMFs as a potential way to reconcile seemingly contradictory results in the literature.